OpenStudy (anonymous):

Find the length of the curve y=x^5/6+ 1/10x^3 from 1< or equal to x < or equal to 2 Arc length equals?

OpenStudy (amistre64):

integrate the length of a hypotenuse

OpenStudy (anonymous):

huh? should i be using the arc length formula? What should my final answer be?

OpenStudy (amistre64):



OpenStudy (anonymous):

?

OpenStudy (amistre64):

as all these terms go infinitesimal we get:\[ds^2=dx^2+dy^2\] \[ds = \sqrt{dx^2+dy^2}\] when y is a function of x, this turns out to be:\[ds = \sqrt{1+(y')^2}\]

OpenStudy (amistre64):

the length of the curve is then the sum of all the itty bitty ds parts:\[\int_C\ ds=\int_{a}^{b}\sqrt{1+(y')^2}\ dx\]

OpenStudy (amistre64):

so what we need to find is y' in order to plug it in what have you got for y' ?

OpenStudy (anonymous):

how would I compute the derivative of y? 5/6+1 bring the 5/6 up front?

OpenStudy (amistre64):

if 5/6 is the exponent then your almost got it :) 5/6 up front and 5/6-1 left

OpenStudy (anonymous):

and then for +1/10x^3 i bring the 3 up front and subtract 1 from exponent, so 3/10x^2?

OpenStudy (anonymous):

srry, I'm so confuse. I appreciate you sitting down and working with me, step by step would help.

OpenStudy (amistre64):

\[y'=\frac{5}{6}x^{1/6}+\frac{3}{10}x^2\] is fine

OpenStudy (amistre64):

y'^2 looks like a nightmare to do, doable, but a pain

OpenStudy (anonymous):

srry.

OpenStudy (amistre64):

\[y'^2=\frac{25}{36}x^{2/6}+\frac{9}{100}x^4+\frac{30}{60}x^{2/6}\] \[y'^2=\frac{25}{36}x^{2/6}+\frac{9}{100}x^4+\frac{3}{6}x^{2/6}\] \[y'^2=\frac{25}{36}x^{2/6}+\frac{9}{100}x^4+\frac{18}{36}x^{2/6}\] \[y'^2=\frac{43}{36}x^{1/3}+\frac{9}{100}x^4\] whew!! i hope i kept that straight

OpenStudy (amistre64):

i see one slight error tho; 5/6 - 6/6 = -1/6 so i have to adjust for that mistake

OpenStudy (anonymous):

Thank you so much!! ill use the procedure you put down for my other questions. I just needed someone to help me get started.

OpenStudy (anonymous):

ah, no worries, I can just look where you went wrong. Thanks again.

OpenStudy (amistre64):

good luck :) those fractions scare me at the moment since I got no idea how thats gonna integrate; unless you can do it by computer

OpenStudy (anonymous):

doing the derivative gets tedious at times especially with the whole computation with 1+ and square root- Yuck. Thanks for walking me through the problem!

OpenStudy (anonymous):

yeah, I totally understand. say, how would I do it on a comp?

OpenStudy (amistre64):

id type it up onto wolframalpha.com and see if it can plow through it; im woking that angle at the moment to see what a mess this things gonna be

OpenStudy (anonymous):

oh.

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sqrt%281%2B%28derivative+x%5E%285%2F6%29%2Bx%5E3%2F10%29%5E2%29 and thats just before you have to integrate it ... ewwww

OpenStudy (amistre64):

to be sure: its spose to be:\[y=x^{5/6}+\frac{1}{10}x^3\]?????

OpenStudy (anonymous):

uh its 1/10^2/3 but i guess no difference. Other than that, yes.

OpenStudy (anonymous):

1< than or equal to x < than or equal to 2

OpenStudy (amistre64):

\[y=x^{5/6}+(\frac{1}{10})^{2/3}\] theres no x in the end there right? if so than thats just a constant that disappears in the derivative

OpenStudy (amistre64):

im thinking your best bet might be to say that sin(t) = y' and use the trig substitution for it

OpenStudy (amistre64):

tan(t) that is ; its a 1+d^2 which comforms better to tan

OpenStudy (anonymous):

no limit

OpenStudy (anonymous):

wait, so what is your final answer? Ill try to input it and see if it's correct.